# A Collectivist, Economic Perspective on AI

### v2

@@ 1 Introduction
Before turning to concrete instances of design using these thinking styles, let’s return to the issue of coping with uncertainty. Rather than attempting to delimit the notion of uncertainty with a precise definition, let’s instead be informal and consider the broad problem of making decisions when some of the information that would be useful in making the decision is not available. Different fields have contributed different kinds of algorithms to cope with uncertainty. Briefly, the field of statistics has focused on uncertainty arising from sampling, whereby the data available for algorithmic processing is a subset of all of the data that would be useful in principle to solve some problem. The field of economics has also made use of sampling ideas, but a distinctive focus of economics has been a different source of uncertainty—the information asymmetry that arises when an agent interacts with another agent who possesses private knowledge and who may strategically reveal or conceal aspects of that knowledge in order obtain to some desired outcome. Information asymmetry does not tend to go away when the sample size grows. Finally, another source of uncertainty comes from the “when, where, and who” of data collection. When data is collected in the past, or at a distant location, or from a different source than the current source, the uncertainty should grow. The field of computer science has contributed tools to address this problem via the algorithmic concept of provenance, whereby the origin and type of data are tracked systematically and coherently. This gives substance to notions of relevance in decision-making that are often treated informally in other fields.
@@ 3 Inference and Incentives
Microeconomics focuses on the choices of strategic agents that are pursuing goals that may be entirely personal. There is also a focus on the overall social welfare that can be achieved when self-interested agents interact. Thus the issues that drive algorithm design in economics include information asymmetries, incentives, social goals, and solutions expressed as equilibria rather than optima.
@@
The theory of incentives builds on game theory, which is a mathematical description of strategic behavior (Roger B, 1991). Indeed, the field of mechanism design, which encompasses the study of incentives, can be viewed as the inverse of game theory: Whereas game theory aims to predict the outcome when strategic agents interact—with the outcome expressed as various kinds of equilibria (e.g., Nash equilibria for simultaneous play and Stackelberg equilibria for sequential play)—mechanism design starts with a desired outcome and asks what game would deliver that outcome as an equilibrium.
@@
One mechanism that is appropriate for sequential play is known as a contract (Laffont and Martimort, 2002). Briefly, the idea is that the Leader does not simply play a single action (e.g., offer a price for some good), but rather presents a menu of options to the Follower, which consists of a set of services and prices. The Follower uses their private knowledge to pick the best option for themselves, and if the Leader has designed the menu well, then many Followers will find an appealing option in the menu. Compared to a mechanism in which a single fixed price is chosen, a contract can deliver higher revenue given that some Followers may opt for a relatively high price in exchange for the services being offered (those Followers with a high “willingness-to-pay”). Moreover, contracts can deliver high social welfare (the aggregrate of the difference between willingness-to-pay and the actual amount paid). Success in achieving such a criterion is expressed in terms of an equilibrium concept; namely, the Stackelberg equilibrium. Algorithm design involves specifying mechanisms (typically decentralized) that yield good Stackelberg equilibria.
@@ 4.2 Data markets
Thus, the user is likely to walk away, and the design needs to be elaborated for the market to function. Let us imagine, for example, that platforms decide to provide a formal guarantee of privacy when sending data to third-party data buyers. Concretely, this would mean that noise is added to the data at a level that is contractually specified (and can be audited). Although the noise level could be subject to government regulation, let’s instead leave the choice in the hands of the platforms. A platform is incentivized to provide a nontrivial level of noise, because users will shop among platforms to find one that provides a desirable level of privacy in conjunction with an effective service. Moreover, such a platform, by accruing more users, will collect more data and can make further improvements to its service. On the other hand, data buyers are averse to noise and will presumably pay less for data from the platforms that provide a stronger privacy guarantee. Thus, there is a conflict for the platforms. The way to understand how the conflict will play out is to model the overall system as a game (it is a generalized Stackelberg game) and find its equilibria. This requires specifying utility functions for the various players.
@@ 4.3 Foundation models, bias, and local knowledge
Foundation models are large-scale ML systems that aim to make high-quality predictions in domains of major scientific or societal interest. For example, LLMs are foundation models for natural language and AlphaFold is a foundation model for protein structure (et al., 2019).
@@ Appendix A Inference and Privacy in Databases
Note that the diagram is not meant to suggest that the randomized algorithm $Q$, which concretely could be a differentially private operator, needs to involve transforming the entire database. Rather, using standard ideas in the differential privacy and online learning literatures (cf. Dwork and Roth, 2014), we envision $Q$ as being called on an as-needed, possibly query-dependent basis.
@@ Appendix B The Missing Middle Kingdom in AI Education
Clearly there is a lot of inferential reasoning going on in this example, and it can be blended with computational concepts in ways that go beyond just plotting histograms. For example, how does one permute a column of numbers in Python? There is a naive $O(n^{2})$ algorithm to do it, but there is also a more sophisticated $O(n)$ algorithm. Aha! Then, critically, with this blend in hand, we returned to the students and asked them to find their own problems, and accompanying datasets, where “treatment” and “control” were whatever real-world distinction they want to investigate. (E.g., does water quality vary between more wealthy parts of a city and well wealthy parts? or Was there more deforestation in the Amazon this past year as compared to five years ago?). Such real-world investigation was exciting for a diverse set of students, in ways that the problems studied in classical computer science curricula are sometimes not (e.g., game-playing programs). Data 8 became the fastest-growing class in the history of UC Berkeley, and is currently taken by over 1500 students each semester.
@@
A related virtue of the tripartite blend is that it is inclusive of the wide range of disciplines that have already developed connections to computer science, economics, and statistics. Indeed, from these three starting points, one arrives quickly at disciplines such as cognitive science, social psychology, law, sociology, public policy, mathematics, physics, biology, and the humanities. These are all disciplines that can provide additional perspectives that are essential in shaping human-centric technology. To take but one example where these perspectives link to our discussion, consider behavioral economics, which links cognitive and social psychology with economics, aiming to overcome a perceived overreliance on unrealistic assumptions of rationality in classical economics. Here too, however, existing concepts need to be revisited in the data-centric era. Recalling that humans provide the data on which AI artifacts such as LLMs are based, we see that behavioral data is in some sense already baked into learning-based systems. Blending the massive-but-messy information in such data with the more traditional, experiment-driven insights coming from behavioral economics is an important agenda item for future research.

### v3

@@ 1 Introduction
Before turning to concrete instances of design using these thinking styles, let’s return to the issue of coping with uncertainty. Rather than attempting to delimit the notion of uncertainty with a precise definition, let’s instead be informal and consider the broad problem of making decisions when some of the information that would be useful in making the decision is not available. Different fields have contributed different kinds of algorithms to cope with uncertainty. Briefly, the field of statistics has focused on uncertainty arising from sampling, whereby the data available for algorithmic processing is a subset of all of the data that would be useful in principle to solve some problem. The field of economics has also made use of sampling ideas, but a distinctive focus of economics has been a different source of uncertainty—the information asymmetry that arises when an agent interacts with another agent who possesses private knowledge and who may strategically reveal or conceal aspects of that knowledge in order to obtain some desired outcome. Information asymmetry does not tend to go away when the sample size grows. Finally, another source of uncertainty comes from the “when, where, and who” of data collection. When data is collected in the past, or at a distant location, or from a different source than the current source, the uncertainty should grow. The field of computer science has contributed tools to address this problem via the algorithmic concept of provenance, whereby the origin and type of data are tracked systematically and coherently. This gives substance to notions of relevance in decision-making that are often treated informally in other fields.
@@ 3 Inference and Incentives
Microeconomics focuses on the choices of strategic agents who are pursuing goals that may be entirely personal. There is also a focus on the overall social welfare that can be achieved when self-interested agents interact. Thus the issues that drive algorithm design in economics include information asymmetries, incentives, social goals, and solutions expressed as equilibria rather than optima.
@@
The theory of incentives builds on game theory, which is a mathematical description of strategic behavior (Myerson, 1991). Indeed, the field of mechanism design, which encompasses the study of incentives, can be viewed as the inverse of game theory: Whereas game theory aims to predict the outcome when strategic agents interact—with the outcome expressed as various kinds of equilibria (e.g., Nash equilibria for simultaneous play and Stackelberg equilibria for sequential play)—mechanism design starts with a desired outcome and asks what game would deliver that outcome as an equilibrium.
@@
One mechanism that is appropriate for sequential play is known as a contract (Laffont and Martimort, 2002). Briefly, the idea is that the Leader does not simply play a single action (e.g., offer a price for some good), but rather presents a menu of options to the Follower which consists of a set of services and prices. The Follower uses their private knowledge to pick the best option for themselves, and if the Leader has designed the menu well, then many Followers will find an appealing option in the menu. Compared to a mechanism in which a single fixed price is chosen, a contract can deliver higher revenue given that some Followers may opt for a relatively high price in exchange for the services being offered (those Followers with a high “willingness-to-pay”). Moreover, contracts can deliver high social welfare (the aggregrate of the difference between willingness-to-pay and the actual amount paid). Success in achieving such a criterion is expressed in terms of an equilibrium concept; namely, the Stackelberg equilibrium. Algorithm design involves specifying mechanisms (typically decentralized) that yield good Stackelberg equilibria.
@@ 4.2 Data markets
Thus, the user is likely to walk away, and the design needs to be elaborated for the market to function. Let us imagine, for example, that platforms decide to provide a formal guarantee of privacy when sending data to third-party data buyers. Concretely, this would mean that noise is added to the data of a magnitude that is contractually specified (and can be audited). Although the noise level could be subject to government regulation, let’s instead leave the choice in the hands of the platforms. A platform is incentivized to provide a nontrivial level of noise, because users will shop among platforms to find one that provides a desirable level of privacy in conjunction with an effective service. Moreover, such a platform, by accruing more users, will collect more data and can make further improvements to its service. On the other hand, data buyers are averse to noise and will presumably pay less for data from the platforms that provide a stronger privacy guarantee. Thus, there is a conflict for the platforms. The way to understand how the conflict will play out is to model the overall system as a game (it is a generalized Stackelberg game) and find its equilibria. This requires specifying utility functions or preferences for the various players.
@@ 4.3 Foundation models, bias, and local knowledge
Foundation models are large-scale ML systems that aim to make high-quality predictions in domains of major scientific or societal interest. For example, LLMs are foundation models for natural language and AlphaFold is a foundation model for protein structure (Jumper et al., 2019).
@@ 5 Discussion
6 Acknowledgments
I would like to acknowledge helpful discussions with Anastasios Angelopoulos, Francis Bach, Stephen Bates, David Blei, Alireza Fallah, Nika Haghtalab, Guido Imbens, Meena Jagadeesan, Barbara Rosario, Ion Stoica, Steve Stoute, Hal Varian, Rakesh Vohra, Serena Wang, and Tijana Zrnic. This work was funded by the European Union, ERC-2022-SYG-OCEAN-101071601. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. I also wish to acknowledge funding by the Chair “Markets and Learning,” supported by Air Liquide, BNP PARIBAS ASSET MANAGEMENT Europe, EDF, Orange and SNCF, sponsors of the Inria Foundation.
@@ Appendix A Inference and Privacy in Databases
Note that the diagram is not meant to suggest that the randomized algorithm $Q$, which concretely could be a differentially private operator, needs to involve transforming the entire database. Rather, using standard ideas from the differential privacy and online learning literatures (cf. Dwork and Roth, 2014), we envision $Q$ as being called on an as-needed, possibly query-dependent basis.
@@ Appendix B The Missing Middle Kingdom in AI Education
Clearly there is a lot of inferential reasoning going on in this example, and it can be blended with computational concepts in ways that go beyond just plotting histograms. For example, how does one permute a column of numbers in Python? There is a naive $O(n^{2})$ algorithm to do it, but there is also a more sophisticated $O(n)$ algorithm. Aha! Then, critically, with this blend in hand, we returned to the students and asked them to find their own problems, and accompanying datasets, where “treatment” and “control” were whatever real-world distinction they want to investigate. (E.g., does water quality vary between more wealthy parts of a city and less wealthy parts? or Was there more deforestation in the Amazon this past year as compared to five years ago?). Such real-world investigation was exciting for a diverse set of students, in ways that the problems studied in classical computer science curricula are sometimes not (e.g., game-playing programs). Data 8 became the fastest-growing class in the history of UC Berkeley, and is currently taken by over 1500 students each semester.
@@
A related virtue of the tripartite blend is that it is inclusive of a wide range of disciplines that have already developed connections to computer science, economics, and statistics. Indeed, from these three starting points, one arrives quickly at disciplines such as cognitive science, social psychology, law, sociology, public policy, mathematics, physics, biology, and the humanities. These are all disciplines that can provide additional perspectives that are essential in shaping human-centric technology. To take but one example where these perspectives link to our discussion, consider behavioral economics, which links cognitive and social psychology with economics, aiming to overcome a perceived overreliance on unrealistic assumptions of rationality in classical economics. Here too, however, existing concepts need to be revisited in the data-centric era. Recalling that humans provide the data on which AI artifacts such as LLMs are based, we see that behavioral data is in some sense already baked into learning-based systems. Blending the massive-but-messy information in such data with the more traditional, experiment-driven insights coming from behavioral economics is an important agenda item for future research.